1. two parallel lines touch the circle at points A and B respectively. if area of the circle is 25pie cm2. then AB is equal to?
2. Area of the greatest circle of the circles constructed on the side of a right triangle, taking these sides as diameters, if the two smalles sides are 1.5cm and 2cm.

Answer by mananth(16949)

Area = 25 pi cm^2
AB are on lines parallel to each other
so AB is the diameter
Area = pi*r^2
25 pi = pir^2
r^2=25
r= 5 cm
AB = 10 cm

-------------
Hypotenuse ^2= leg1^2+leg2^2

leg1 = 1.5 cm
Leg2^2= 2 cm

Hypotenuse^2 1.5 ^2 + 2 ^2
Hypotenuse^2 2.25 + 4
Hypotenuse^2 6.25
take the square root
Hypotenuse^2 2.5 cm
the diameter of the circle = 2.5 cm
radius = 1.25
Area of circle = pi*r^2
Area =pi*1.25^2
Area = 4.90 cm^2
----------